i've been stuck on this all night and cant figure it out (for a homework) and i need to know the answer and how i got to it: 6 tractors take 11 days to collect the harvest. how long would it take 15 tractors to do the same work?
1
Expert's answer
2020-04-27T17:57:32-0400

The amount of the work is "11*6=66" tractor*days or "15*x" Tractor*days.

So,

"66=x*15"

"x=\frac{66}{15}=4.4" days.

It will take 4.4 days for 15 tractors to do the work